OpenStudy (mandre):

How do you calculate the derivative of an absolute value, e.g. g' if g(x) = |x - 2|

OpenStudy (amistre64):

its similar to calculating: sqrt(x^2)

terenzreignz (terenzreignz):

It all boils down to the fact that absolute value is basically just a piecewise function in disguise. In particular \[\Large g(x) = |x-2|= \left\{\begin{matrix}x-2 &&& x\ge2\\\\2-x&&&x<2\end{matrix}\right.\]

OpenStudy (mandre):

I saw that (square root of a square) in a Google search but my study guide shows something else. I think I'll ask this again once I get home to check my study guide.

OpenStudy (amistre64):

\[\sqrt{x^2}~:~let~u=x^2\] \[\sqrt{u}\to\frac{u'}{2\sqrt{u}}\to\frac{2x}{2\sqrt{x^2}}\to\frac{x}{|x|}\]

terenzreignz (terenzreignz):

oh....^ Das geht... lol

OpenStudy (amistre64):

:)

OpenStudy (anonymous):

@terenzreignz Love the way you explained that. It actually makes learning absolute value more fun. "...in disguise..". lol.

terenzreignz (terenzreignz):

Yeah... most unfortunately, it'd probably be a headache if it were more than just the absolute value of some linear function... imagine polynomials... we'd have to find its roots D:

OpenStudy (anonymous):

hehehehehe @terenzreignz But if you know the basic definition, no matter how difficult the problem, you'd be able to figure it out =]

OpenStudy (mandre):

Any time now someone can show me how to do a derivate of |x-2| lol. Bear in mind that I only started calculus a few weeks ago and we didn't have calculus at high school.

OpenStudy (anonymous):

@Mandre Now according to the piecewise function @terenzreignz gave you, differentiate each and see what you get. Each derivative will correspond to the respective interval.

terenzreignz (terenzreignz):

And you'll see that it jives (fortunately) with amistre's more elegant derivative :)

terenzreignz (terenzreignz):

\[\Large \frac{d}{dx}g(x) = \frac{d}{dx}|x-2|= \left\{\begin{matrix}\frac{d}{dx}[x-2] &&& x\color{red}{>}2\\\\\\\\\frac{d}{dx}[2-x]&&&x<2\end{matrix}\right.\]

OpenStudy (mandre):

Then it's 1 for (x-2) or -1 for (2-x)?

OpenStudy (mandre):

why not? Something to do with continuity?

terenzreignz (terenzreignz):

@Mandre That is correct :) Notice I replaced the \(\ge\) sign in the top-function of the piecewise with a simple > sign... That is because the derivative simply doesn't exist at the point x = 2.

OpenStudy (mandre):

I need to revise my continuity. It seemed so simple at the time * sigh *

OpenStudy (amistre64):

my derivative is undefined for x=2 ...

OpenStudy (anonymous):

@amistre64 how is it undefined at '2'?

terenzreignz (terenzreignz):

@amistre64 \[\Large \frac{x-2}{|x-2|}\]?

OpenStudy (mandre):

Div by 0

OpenStudy (amistre64):

\[\frac d{dx}|x-2|\to\frac{x-2}{|x-2|}\]

OpenStudy (amistre64):

since |x-2| has no derivative at x=2 to begin with ....

OpenStudy (anonymous):

Oh it's (x-2). Sorry I forgot that lol =P

OpenStudy (amistre64):

my initial post was really about a linear absolute value in general

OpenStudy (anonymous):

@amistre64 oh ok lol. I thought that it was the question rofl...

OpenStudy (amistre64):

i had made an explicit rule for some sequence in class .... using the derivative of a combination of linear absolute values such that at any given "n", it matched the stated sequence .... over kill, but it helped to stem the tide of boredom

OpenStudy (mandre):

Here is my actual problem. Find the derivative of g \[g(z) = \frac{ z ^{3}-z+1 }{|1-2z| }\]

OpenStudy (mandre):

I know how to use the quotient rule but get unstuck at the abs value.

OpenStudy (amistre64):

you could either divvy it up into parts like terens did, or use my demonstration

terenzreignz (terenzreignz):

Oh, lol... we can't have piecewise functions, now can we? :P You better use amistre's derivative...

OpenStudy (amistre64):

\[g(z) = \frac{ z ^{3}-z+1 }{(1-2z) };~\frac{ z ^{3}-z+1 }{(2z-1) }\]

OpenStudy (mandre):

Thank you amistre. You're a star.

OpenStudy (mandre):

Thanks everyone

OpenStudy (amistre64):

yep, that me ... the lorenzo lamas of openstudy lol

OpenStudy (mandre):

lol.